OpenStudy (anonymous):

√a^7√6a^3 Multiply and simplify, if possible. The variable represents a nonnegative real number.

OpenStudy (lalaly):

\[\sqrt{a^7} \sqrt{6a^3} = a^{\frac{7}{2}} (6a^3)^{\frac{1}{2}}\]\[=\large{\sqrt6 a^{\frac{7}{2}+\frac{3}{2}}}\]

OpenStudy (lalaly):

\[=\sqrt6 \times a^{\frac{10}{2}} = \sqrt6 \times a^5\]

jhonyy9 (jhonyy9):

\[=a ^{5}\sqrt{6}\]